OpenStudy (anonymous):

Ugh, I cant do math for my life. Evaluate the expression for m= -5 n= 4/3 p= -9 -3mnp

OpenStudy (mathstudent55):

Can you multiply numbers together?

OpenStudy (mathstudent55):

Replace m, n, and p with the values you are given. -3mnp = -3(-5)(4/3)(-9) Now multiply the numbers together. Start with (-3)(-5)

OpenStudy (anonymous):

Yes cx @mathstudent55

OpenStudy (igreen):

-3mnp Input the given numbers: -3(-5)(4/3)(-9) Multiply from left to right: 15(4/3)(-9) 20(-9) -180

OpenStudy (bibby):

Stop spoonfeeding @iGreen

OpenStudy (anonymous):

-3 * -5 = 15. @mathstudent55 Oh and thanks again xD @iGreen

OpenStudy (mathstudent55):

Great. Now you do the next step: -3mnp = -3(-5)(4/3)(-9) = 15(4/3)(9) Now what is 15 * 4/3 = ?

OpenStudy (anonymous):

11.25

OpenStudy (anonymous):

1/15 * 4/3 ( I think is correct...?) and then you get 4/15....?

OpenStudy (mathstudent55):

Sorry, I left he negative out of the -9. I did it again below: Great. Now you do the next step: -3mnp = -3(-5)(4/3)(-9) = 15(4/3)(-9) Now what is 15 * 4/3 = ?

OpenStudy (mathstudent55):